OpenStudy (anonymous):

What is the exact value of the expressions sqrt 180 − sqrt 125 + sqrt 5 ? Simplify if possible.

OpenStudy (anonymous):

\[\sqrt{180}- \sqrt{125} + \sqrt{5}\]

OpenStudy (ghazi):

\[\sqrt{180}= 6\sqrt{5}....\sqrt{125}= 5\sqrt{5}....\sqrt{5}\] hope you can do it now

OpenStudy (ghazi):

can you solve this....?

OpenStudy (anonymous):

not really. i dont know how to do this? can you explain?

OpenStudy (ghazi):

it's just simplification ....\[6\sqrt{5}-5\sqrt{5}+\sqrt{5}= 2\sqrt{5}\]

OpenStudy (ghazi):

how i got these ...check out the comments

OpenStudy (anonymous):

okay thanks

OpenStudy (ghazi):

:)